# Spatial Scale of Agglomeration and Dispersion Number, Spacing, and the Spatial Extent of Cities

### v7

@@ 3.4   Evolution of spatial patterns
Another important caveat about Proposition 1 is that it relies on local stability analysis at $\bar{\bm{x}}$, and hence does not tell whether the instability of $\bar{\bm{x}}$ toward a polycentric deviation pattern indeed implies the existence of stable polycentric spatial equilibria in the vicinity of $\phi^{*}$. In fact, to study the path of stable spatial equilibria along the transport cost axis beyond Proposition 1, one must assume a specific indirect utility function, and thus possible characterizations are inherently model dependent. [^18] For example, Kucheryavyy-etal-JIE2024 focused on a specific but flexible two-region model that encompass Allen-Arkolakis-QJE2014 and Krugman-JPE1991 as special cases. They essentially showed that the agricultural sector à la Krugman-JPE1991 produces a global dispersion force that stabilize symmetry at high transport costs, which is consistent with our results. On the other hand, existence results for equilibria are available in general symmetric geographies (Ikeda-Murota-Book2014).
@@ B The Forslid–Ottaviano model (Type G)
Figure 8: Numerical illustration of Propositions 3 and 2
Nonetheless, there is indeed a sharp difference in multi-region agglomeration process under the influence of different dispersion mechanisms. In particular, polycentric spatial patterns can become stable only under a pronounced global dispersion force, which is consistent with Proposition 1. This is illustrated by the following formal results for selected models from Types L and G under Assumption C with $N=4$:
Proposition 2.
Consider the Type L model by Helpman-Book1998 in the four-region symmetric circle. Assume that multiple equilibria exists, and that the full dispersion $\bar{\bm{x}}=(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$ is stable for large $\phi$. Then, there exists $\phi^{*}\in(0,1)$ such that $\bar{\bm{x}}$ is unstable for all $\phi\in(0,\phi^{*})$. Duocentric spatial equilibria of the form $(m_{1},m_{2},m_{1},m_{2})$ with $m_{1}>m_{2}$, if exist, are unstable for all $\phi\in(0,\phi^{*})$, implying that stable equilibria must be single-peaked for this range. The full dispersion $\bar{\bm{x}}$ is stable for all $\phi\in(\phi^{*},1)$, and a monocentric equilibrium path is connected to $\bar{\bm{x}}$ at $\phi^{*}$ [Proposition 1 (a)].
Proposition 3.
Consider the Type G model by Forslid-Ottaviano-JoEG2003 in the four-region symmetric circle. Assume that the full dispersion $\bar{\bm{x}}=(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$ is stable for small $\phi$. Suppose that the initial state is $\bar{\bm{x}}$, and $\phi$ increases monotonically from 00. At some $\phi^{*}\in(0,1)$, $\bar{\bm{x}}$ becomes unstable as the duocentric deviation pattern $\bm{z}=(1,-1,1,-1)$ generates positive utility gain [Proposition 1 (b)]. The new stable equilibrium is the duocentric concentration, e.g., $(\frac{1}{2},0,\frac{1}{2},0)$. At some $\phi^{**}>\phi^{*}$, the duocentric equilibrium becomes unstable, as the deviation pattern $\bm{z}=(1,0,-1,0)$ now generates a positive utility gain. Finally, the full concentration in a single region such as $(1,0,0,0)$ becomes the stable spatial equilibrium for large $\phi$.
Proof of Propositions 3 and 2.
See AMT-2016. ∎
Figure 8 numerically illustrates Propositions 3 and 2, with schematic diagrams above showing the corresponding spatial patterns along the circular space. In particular, in the Helpman model, a monocentric distribution of the form $\bm{x}=(m_{2},m_{1},m_{2},m_{3})$ with $m_{1}>m_{2}>m_{3}$ is the stable equilibrium for $\phi\in(0,\phi^{*})$. As $\phi$ increases, it converges to the full dispersion, which is consistent with Proposition 1 (a).
As a further numerical example, Figure 9 depicts the typical evolution of spatial patterns, assuming $N=16$. We consider a monotonic decrease in transport costs (a monotonic increase in $\phi$), and follow a path of stable spatial equilibria.
@@ C Type LG
Note: For interpretability, the spatial distributions in the circular economy are visualized as if it is on a line segment. In each figure, the leftmost region is neighboring to the rightmost one under Assumption C.
@@
Figure 9A considers the Type L model by Helpman-Book1998. The symmetry $\bar{\bm{x}}$ is unstable if $\phi$ is small, and agents concentrate around a single peak. As $\phi$ increases, it causes the spreading of the single agglomeration as the peak population density is reduced. At a critical level of $\phi$, the spatial distribution converges to $\bar{\bm{x}}$ [Proposition 1 (a)]. In particular, stable equilibria are single-peaked throughout the process.
@@
Figure 9B considers Krugman-JPE1991’s model as a leading example of Type G. At low levels of $\phi$, the uniform distribution is stable. As we increase $\phi$, an endogenous transition occurs at a critical threshold of $\phi$ [Proposition 1 (b)]: a polycentric pattern with eight agglomerations becomes a stable equilibria. This eight-centered configuration remains stable over a range of $\phi$. Further increases in $\phi$ reduce the number of agglomerations, widen the spacing between them, and enlarge each center, via successive instability of equilibria similar to Proposition 3. The number of economic centers in stable spatial equilibrium evolves as $16\to 8\to 4\to 2\to 1$, reflecting fewer and more widely spaced agglomerations as transport costs fall. [^19] In this class of models, established centers may decline because the growth of larger agglomerations comes at the expense of smaller ones. For example, in Fig. 9B, the fifth region from the left initially gains population as transport improves but later loses population due to increasing concentration elsewhere.
@@
Type LG combines the properties of Types L and G. Figure 9C considers a Type LG model proposed by Pfluger-Suedekum-JUE2008. The uniform distribution is stable when $\phi$ is small. As $\phi$ increases, octacentric agglomerations emerge at some point as in Type G models. Multiple bell-shaped economic agglomerations are generated at moderate levels of transport costs. Further increase in $\phi$ causes a decrease in the number of agglomerations (as in Type G) and the spreading of each agglomeration (as in Type L). When $\phi$ is close to one, the spatial distribution becomes monocentric, similar to Type L. At the low transport cost regime, in particular, the model exhibits a transformation from a two-peaked to a single-peaked pattern accompanied by local spreading, which resembles the dual evolution of cities, i.e., global concentration and local flattening, discussed in Section 1. While the real-world dynamics appear to unfold simultaneously, the overall evolutionary process in the model generates these changes somewhat sequentially: first through a reduction in the number of peaks, and then through the flattening of the remaining agglomeration. This discrepancy may reflect limitations of the model, the most fundamental of which would be its static structure.
4   Concluding discussions
This study illustrated how the spatial scale—local or global—of dispersion forces embedded in a many-region spatial model may lead to qualitatively different implications on equilibrium spatial patterns of economic activities and their comparative statics. Leveraging the tractability of the symmetric circular geography, this study provides a conceptual lens through which the distinction between local and global forces can be formalized, offering a general framework for interpreting results across theoretical, quantitative, and empirical research. The robustness of our theoretical conclusions beyond the circular economy is a crucial issue. Below, we provide brief explorations on this respect with discussion on related literature, leaving more thorough and rigorous analysis for future research.
@@ B The Krugman model (Type G)
Figure 10: Stable spatial patterns in a square economy ($819^{2}=$ 81 locations).
Geographical advantages. The insights from the circular geography appear to qualitatively generalize to other setups, in which the shape of network produces innate geographical advantages. For illustration, Figure 10 compares stable equilibria for models from Types L and G in a square geography for different level of transport costs, where local regional fundamentals are homogeneous. In this geography, central regions have better accessibility compared with the peripheral regions. Similarities between Fig. 10 and Figs. 9A and 9B are evident. In particular, the spatial distribution of agents become monocentric in Type L, whereas it becomes polycentric in Type G. See Appendix D for discussions. To obtain formal results in such geographical settings, a feasible strategy would be to select a few tractable but representative models for each model type and then identify general insights when the network structure systematically varies, as Matsuyama-RIE2017 did in the context of a trade model.
Local advantages. Instead of incorporating advantages that stem from the underlying network structure, one can introduce regional heterogeneity while retaining the circular network. Appendix E formally examines how small differences in intrinsic local characteristics, such as productivity or amenities, affect the spatial equilibrium and its sensitivity to interregional transport costs. Specifically, it shows that the model’s internal dispersion mechanism plays a crucial role in determining whether the effects of such local advantages are amplified or mitigated as transport accessibility improves.
@@
The magnitude of $\rho$ at a given transport cost level captures how the model’s endogenous forces translate variation in $\bm{a}$ into variation in regional population distribution. The key insight from Appendix E (see Propositions 4 and 5) is that the direction of $\rho$’s response to changes in transport costs is largely determined by the type of dispersion force embedded in the model. In models with global dispersion mechanisms, improved interregional access tends to magnify initial local advantages (i.e., $\rho$ increases), as population concentrates further in already favored regions. By contrast, in models with local dispersion mechanisms, the same improvement tends to dampen the effects of innate heterogeneity, leading to a more even population distribution (i.e., $\rho$ decreases). These patterns are also consistent with the simple examples shown in Figs. 3D and 3C.
@@ C Removal of highways (Type LG)
Figure 11: Actual and counterfactual population changes of Japanese regions.
@@
Note: The figures are reproduced from Sugimoto-etal-DP2025. Panel (A) shows the actual population changes from 2020 to 1970 (i.e., backward in time). Panel (B) applies a many-region version of Helpman-Book1998 (Type L). In Panel (C), Sugimoto-etal-DP2025 proposes a Type LG model by extending the Helpman framework by incorporating land (as in Pfluger-Tabuchi-RSUE2010) and intermediate goods (as in Krugman-Venables-EER1996) as additional inputs in production. Importantly, the Type LG model does not assume immobile workers. Instead, land as the immobile factor produces a global dispersion force, as in Pfluger-Tabuchi-RSUE2010. For further details and discussion, see Sugimoto-etal-DP2025.
Geographical and local advantages. The type of dispersion mechanism can lead to markedly different comparative statics even in real-world geographies with both geographical and local heterogeneity. Figure 11 illustrates observed and counterfactual population changes across Japanese regions, focusing on the long-run impact of highways. Panel (A) shows actual changes from 2020 to 1970 (i.e., backward in time), highlighting the historical shift toward central regions over the past 50 years. Panels (B) and (C) present counterfactual simulations using the calibrated Type L and Type LG models, respectively. Each model is calibrated to the 2020 population distribution and used to assess the effect of removing the highway network (Fig. 1A).
While neither model is expected to replicate the actual evolution of population patterns, a striking contrast emerges between the two. In the Type L model, which features only local dispersion forces, the removal of highways leads to further concentration in core regions such as Tokyo. This behavior—greater centralization as transport costs rise—is consistent with the theoretical properties of Type L models. By contrast, the Type LG model, which incorporates both local and global dispersion forces, produces the opposite outcome: the same highway removal triggers a substantial population shift toward peripheral regions. This qualitative reversal highlights how the nature of embedded dispersion forces can fundamentally shape counterfactual implications in quantitative applications.
@@
These observations call for further research on the role of endogenous economic forces in quantitative spatial models (QSMs). Redding-Rossi-Hansberg-ARE2017 states that the literature “does not aim to provide a fundamental explanation for the agglomeration of economic activity, but rather to provide an empirically relevant quantitative model to perform general equilibrium counterfactual policy exercises” (p.23). Building on this premise, the literature interprets unexplained interregional variation in endogenous outputs of the model (i.e., structural residuals) as innate location fundamentals. Hence, “agglomeration in these models is simply the result of exogenous local characteristics augmented by endogenous economic mechanisms” (Ibid.), and endogenous forces have played less featured roles than they did in stylized theories of spatial agglomeration. [^20] In Redding-Sturm-AER2008 and Allen-Arkolakis-QJE2014, the replication data indicate that the structural residuals (interpreted as housing stock in the former, and unobserved amenity and productivity in the latter) account for 90 % and 78 % of the log-city size variation in their context. However, as demonstrated by the simple experiment in Fig. 11, endogenous dispersion mechanisms can have a significant qualitative impact on counterfactual results in QSMs and hence policy suggestions.
Baum-Snow-etal-JUE2020 notes that, in the context of Chinese urban system, “the presence of an urban hierarchy appears to require either returns to scale that are important enough to permit multiple equilibria as in Krugman-JPE1991, or else an important role for industrial specialization that depends on prefecture abundance of land or natural resources,” highlighting the possible importance of strong endogenous forces as well as immobile factors. That said, an exciting yet challenging quantitative question for future research would be whether placing greater emphasis on endogenous economic forces enhances the explanatory power and/or empirical fit of QSMs.
Generally, spatial models that emphasize only one type of dispersion force might capture only certain aspects of how economic activity responds to policy and technological changes. Notably, most regional-scale quantitative models emphasize local dispersion force as the spreading mechanism, and thus tend to predict population decentralization in response to reduced transport costs, as Fig. 11B illustrates. Our results suggest that, to capture the differential regional impacts—the winners and losers—due to new large-scale transport infrastructures, accounting for global dispersion forces can be crucial, warranting further theoretical exploration of their microfoundations. See also Graham-Horcher-RP2024 for a thoughtful evaluation of conventional QSMs as potential tools for ex-ante welfare analysis of transport interventions in practice.
Empirical connections. Notably, our framework provides a unifying and coherent interpretation of various empirical findings in the literature. For example, the evidence on regional agglomeration presented by Duranton-Turner-REStud2012 and Faber-REStud2014 aligns with the predictions of Type G models. The former focused on the growth of large metropolitan areas in the United States, while the latter examined the growth of peripheral counties in China. The former study revealed a positive correlation between regional growth rates and the level of transport infrastructure, whereas the latter found a negative correlation. These opposite responses can be interpreted as different sides of the same coin for Type G models: both results indicate a tendency of concentration towards bigger regions when interregional transport access improves. Baum-Snow-etal-JUE2020 also provided an evidence for global concentration across Chinese prefectures due to the introduction of the Chinese national highway system by showing that hinterland prefectures experienced slower growth rates compared to regional primates.
Internal structure of cities and intercity distribution of economic activities are often discussed separately, and an interesting subtlety emerges here. The number of interregional highway linkages (e.g., the number and total length) within a given region is often used as a measure of intra-urban transport infrastructure, including Duranton-Turner-REStud2012 discussed above. Here, we interpret this variable instead as an indicator of exposure to improvements in the inter-regional transport network. This interpretation allows an alternative explanation of their results in the context of Type G models. This highlights the importance of a unified analytical perspective that bridges interregional and intraurban spatial structures, where the same empirical measure may carry distinct theoretical meanings depending on the spatial scale of interpretation.

### v8

@@ 3.4  Evolution of spatial patterns
Because Proposition˜1 relies on a local stability analysis around $\bar{\bm{x}}$, it does not establish whether an instability toward polycentric deviations actually leads to stable polycentric equilibria. To trace how stable spatial equilibria evolve as transport costs change beyond Proposition˜1, one must specify an indirect utility function, and the resulting characterizations are therefore model dependent. [^19] For example, Kucheryavyy-etal-JIE2024 focused on a specific but flexible two-region model that encompasses Allen-Arkolakis-QJE2014 and Krugman-JPE1991 as special cases. They essentially showed that the agricultural sector à la Krugman-JPE1991 produces a global dispersion force that stabilizes the symmetric equilibrium at high transport costs, which is consistent with our results.
Nonetheless, in specific models, we can show that polycentric spatial patterns become stable only when global dispersion forces are sufficiently strong, consistent with Proposition˜1. This is illustrated by the following formal results and Fig.˜9.
Proposition 2.
Consider the Type L model of Helpman-Book1998 on a symmetric four-region circle. Suppose that the full dispersion $\bar{\bm{x}}=(\tfrac{1}{4},\tfrac{1}{4},\tfrac{1}{4},\tfrac{1}{4})$ is stable for large $\phi$, but that multiple equilibria may exist. Then there exists a threshold $\phi^{*}\in(0,1)$ such that $\bar{\bm{x}}$ is unstable for all $\phi\in(0,\phi^{*})$. For this range, any duocentric equilibrium of the form $(m_{1},m_{2},m_{1},m_{2})$ with $m_{1}>m_{2}$, if it exists, is also unstable, implying that all stable equilibria must be single-peaked. For $\phi\in(\phi^{*},1)$, the fully dispersed allocation $\bar{\bm{x}}$ is stable, and a single-peaked equilibrium path connects to $\bar{\bm{x}}$ at $\phi^{*}$ [cf. Proposition˜1 (a)].
Proposition 3.
Consider the Type G model by Forslid-Ottaviano-JoEG2003 on a four-region symmetric circle. Assume that the full dispersion $\bar{\bm{x}}=(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$ is stable for small $\phi$. Suppose that the initial state is $\bar{\bm{x}}$, and $\phi$ increases monotonically from 0. At some threshold $\phi^{*}$, $\bar{\bm{x}}$ becomes unstable, and duocentric spatial equilibrium of the form $(m_{1},m_{2},m_{1},m_{2})$ with $m_{1}>m_{2}$ branches from $\bar{\bm{x}}$ [cf. Proposition˜1 (b)]. In particular, the new stable equilibrium is the duocentric agglomeration such as $(\frac{1}{2},0,\frac{1}{2},0)$. At some $\phi^{**}>\phi^{*}$, the duocentric equilibrium becomes unstable. Finally, the full agglomeration in a single region such as $(1,0,0,0)$ becomes the stable spatial equilibrium for large $\phi$.
Proof of Propositions˜3 and 2.
See AMT-2016. ∎
@@ B The Forslid–Ottaviano model (Type G)
Figure 9: Numerical illustration of Propositions˜3 and 2.
In Fig.˜9, the schematic diagrams above show the corresponding spatial patterns. In particular, in the Helpman model, a monocentric distribution of the form $\bm{x}=(m_{2},m_{1},m_{2},m_{3})$ with $m_{1}>m_{2}>m_{3}$ is the stable equilibrium for $\phi\in(0,\phi^{*})$. It converges to the full dispersion at $\phi^{*}$, which is consistent with Proposition˜1 (a).
As a further numerical example, Figure˜10 shows the typical evolution of spatial patterns, assuming $N=16$. We consider a monotonic increase in $\phi$ (i.e., a monotonic decrease in transport costs), and follow a path of stable spatial equilibria.
@@ C Type LG
Note: The spatial distributions in the circular economy are visualized as if it is on a line segment. In each figure, the leftmost region is neighboring to the rightmost one under Assumption˜C.
@@
Figure˜10A considers the Type L model by Helpman-Book1998. The symmetric equilibrium $\bar{\bm{x}}$ is unstable if $\phi$ is small and the agents concentrate around a single peak. As $\phi$ increases, the monotonic spread of the single-peaked distribution occurs. At a critical level of $\phi$, the spatial distribution converges to $\bar{\bm{x}}$ [Proposition˜1 (a)]. In particular, stable equilibria are single-peaked throughout the process.
@@
Figure˜10B considers the Type G model of Krugman-JPE1991. When $\phi$ is low, the symmetric equilibrium is stable. As $\phi$ increases, an endogenous transition occurs, and a polycentric equilibrium with eight agglomerations becomes stable [cf. Proposition˜1 (b)]. Further increase in $\phi$ triggers successive instabilities: the number of agglomerations falls, the spacing between them widens, and each remaining center grows larger (cf. Proposition˜3). As a result, the number of centers in the stable equilibrium evolves as $16\to 8\to 4\to 2\to 1$. In Type G models, spatial adjustment generates both winners and losers. Centers that initially grow may later decline as larger agglomerations expand at their expense. For instance, in Fig.˜10B, the fifth region from the left initially gains population as $\phi$ increases but eventually loses population.
@@
Figure˜10C considers a many-region version of the Type LG model by Pfluger-Suedekum-JUE2008. The symmetric equilibrium is stable if $\phi$ is small. As $\phi$ increases, eight-centric agglomerations emerge at some point, as in Type G. Multiple bell-shaped agglomerations are generated at moderate $\phi$. Increasing $\phi$ further causes a decrease in the number of agglomerations and the spread of each agglomeration. When $\phi$ is close to one, the economy becomes monocentric, as in Type L. Notably, in the large $\phi$ regime, the model exhibits a transformation from a two-peaked to a single-peaked pattern accompanied by local spreading. This broadly resembles the dual evolution of cities discussed in Section˜1.202020While the real-world dynamics appear to unfold simultaneously, the model generates these changes somewhat sequentially: first through a reduction in the number of peaks, and then through the flattening of the remaining agglomeration. This discrepancy may reflect limitations of the model, the most fundamental issues being the absence of inter-location commuting and dynamic decisions.
4  Asymmetries
Real-world geography differs markedly from the stylized benchmarks discussed so far. Geographic accessibility and other region-fixed attributes can vary between locations, generating distortions absent in idealized settings. Yet, as a brief exploration in this section demonstrates, these exogenous asymmetries do not alter the core insights: the spatial scale of dominant dispersion mechanisms can fundamentally shape the spatial distribution and their response to transport costs.
@@ B The Krugman model (Type G)
Figure 11: Stable spatial patterns in a square economy ($9^{2}=81$ locations).
4.1  Geographic accessibility
Asymmetries can arise solely from the underlying distance structure between regions. To illustrate this point, Figure˜11 considers a square geography with homogeneous local fundamentals and compares stable equilibria under the Type L and Type G models. Unlike in a circular economy, some regions are “central” and therefore enjoy an inherent accessibility advantage. Nonetheless, our theoretical result continues to hold: the spatial distribution is monocentric under Type L, whereas it is polycentric under Type G. Moreover, comparative statics with respect to reductions in transport costs are qualitatively similar to those in Figs.˜10A and 10B: under Type L, population spreads monotonically, while under Type G, agglomeration proceeds through successive concentration into fewer locations. See Appendix D for further examples. To obtain formal results for such asymmetric settings, one approach is to select tractable models representative of each type and study their implications across alternative network structures, following the strategy of Matsuyama-RIE2017 in a trade context.
4.2  Local fundamentals
Other than innate accessibility advantages, region-specific parameters such as exogenous productivity and amenity levels are equally fundamental, especially in quantitative spatial models (QSMs). For example, in the model of Allen-Arkolakis-QJE2014, indirect utility can be written as
@@ 4.2  Local fundamentals
where $\{w_{i}\}$ are the nominal wages determined in the market equilibrium given $\bm{x}$:
$w_{i}x_{i}=\sum_{j\in\mathcal{I}}\frac{w_{i}^{1-\sigma}\big(\overbrace{b_{i}x_{i}^{\alpha}}^{\mathclap{\text{Within-region productivity spillover (local agglomeration force)}}}\big)^{\sigma-1}\phi_{ij}}{\sum_{k\in\mathcal{I}}w_{k}^{1-\sigma}\left(b_{k}x_{k}^{\alpha}\right)^{\sigma-1}\phi_{kj}}w_{j}x_{j}$ $\forall i\in\mathcal{I},$ (40)
The key elasticities are $\alpha>0$, $\beta>0$, and $\sigma>1$. The equilibrium is unique if $\alpha-\beta<0$. In this case, given the observed population vector $\hat{\bm{x}}$, we can uniquely solve for the region-specific parameters $u_{i}$} and $b_{i}$} that rationalize $\hat{\bm{x}}$ as the model’s equilibrium. A natural question then is how such regional differences affect our results.
To address this question, the circular economy remains useful. Consider the proximity structure as in Assumption˜C, but allow for variations in region-specific amenities, which we denote by $\bm{a}=(a_{i})_{i\in\mathcal{I}}$ with $a_{i}>0$. When $a_{i}=\bar{a}$ for all $i$, we recover the symmetric racetrack economy of Section˜3, for which $\bar{\bm{x}}$ is an equilibrium. If we slightly perturb $\bm{a}$ from $\bar{\bm{a}}\equiv(\bar{a},\bar{a},\ldots,\bar{a})$, then $\bar{\bm{x}}$ is also slightly perturbed to form a new equilibrium $\bm{x}(\bm{a})$, which can be seen as a function of $\bm{a}$.
To summarize the overall effect of such heterogeneities in $\bm{a}$ on the spatial distribution, we can use the covariance between each region’s relative advantage and its deviation in population share from $\bar{\bm{x}}$:
[…]
If $\rho=0$, heterogeneity in $\bm{a}$ does not affect the spatial distribution. We assume $\rho>0$ without loss of generality, as more advantaged regions should attract more population.
Since the mapping $\bm{x}(\bm{a})$ is model dependent, the magnitude of $\rho$ at a given transport cost level captures how the model’s endogenous forces translate variation in $\bm{a}$ into variation in regional population distribution. Appendix E formally shows that the sensitivity of $\rho$ to transport costs differs markedly between models, depending on the spatial scale of the dispersion forces (Propositions˜4 and 5).
@@ C Removal of highways (Type LG)
Figure 12: Observed and counterfactual population change rates of Japanese regions.
@@
Note: Figures are based on the simulation data of Sugimoto-etal-DP2025. Panel (B) reports a many-region version of the Type L model of Helpman-Book1998. Panel (C) reports the Type LG model of Sugimoto-etal-DP2025, which extends the Helpman framework by incorporating land and intermediate goods as additional inputs. Although Sugimoto et al. does not assume immobile workers, land as an immobile factor generates a global dispersion force, as in Pfluger-Tabuchi-RSUE2010.
Specifically, in models with pronounced global dispersion forces, improved interregional access tends to magnify initial local advantages: $\rho$ increases as the freeness of transport $\phi$ increases and the population becomes more concentrated in the regions favored due to exogenous advantages. By contrast, in models with strong local dispersion forces, the same transport improvement tends to dampen the role of innate heterogeneity, resulting in a flatter distribution, and $\rho$ decreases as $\phi$ increases. These patterns are consistent with the two-region examples in Figs.˜4D and 4C.
4.3  The combination: A quantitative example
In reality, both geographic accessibility and local fundamentals vary between regions. We briefly discuss Sugimoto-etal-DP2025’s results that compare Type L and Type LG over the Japanese geography. Figure˜12 reports the observed and counterfactual population growth rates. Panel (A) shows actual changes from 2020 to 1970 (i.e., backward in time), highlighting the shift to major centers such as Tokyo, Osaka, and Nagoya. Panels (B) and (C) present counterfactual simulations based on calibrated Type L and Type LG models, respectively. Each model is calibrated to the population distribution observed in 2020 under the assumption of a unique equilibrium and then used to evaluate the counterfactual effects of removing the highway network.
A Calibration to the population in 2020
B Counterfactual growths
Figure 13: Comparison of model behaviors in Fig.˜12
Note: Panel (A): In both models, beyond 9090% of log population variation is explained by exogenous region-fixed fundamentals (a composite of observed land area and unobserved amenities). Panel (B): In the Helpman model (Type L), growth rates and exogenous fundamentals are positively associated, i.e., highway removal induces growths at high-amenity regions, which are, as Panel (A) shows, essentially more populated regions in 2020. The converse holds true for Sugimoto et al.’s Type LG model as we observe negative association between exogenous fundamentals and growth rates.
There is a notable contrast between the two models. In the Type L model, the removal of highways leads to further concentration in the core regions (Fig.˜12B). This aligns with the theoretical property of Type L models that higher transport costs induce greater centralization. Figure˜12C shows the opposite pattern: in the Type LG model, the same shock produces a substantial population shift toward peripheral regions.
Figure˜13B uses the same simulation data as Figs.˜12B and 12C to show that, as transport costs rise, more advantaged regions grow faster in Type L, whereas the opposite occurs in Type LG. This pattern is consistent with the discussion in Section˜4.2. These qualitative reversals illustrate that the embedded forces can fundamentally shape the counterfactual implications of spatial models in asymmetric geographies.
4.4  On quantitative spatial models
@@
These results call for research on the role of endogenous economic forces in QSMs. A central premise of the QSM literature is that it “does not aim to provide a fundamental explanation for the agglomeration of economic activity, but rather to provide an empirically relevant quantitative model to perform general equilibrium counterfactual policy exercises” (Redding-Rossi-Hansberg-ARE2017, p. 23). Building on this premise, QSMs typically impose equilibrium uniqueness and interpret unexplained interregional variation (i.e., structural residuals) as innate regional fundamentals. As a result, much of the observed variation is attributed to structural residuals (cf. Fig.˜13A), and endogenous forces play a more limited role than in stylized theories of agglomeration. [^21] For example, in the seminal works of Redding-Sturm-AER2008 and Allen-Arkolakis-QJE2014, structural residuals account for 90% and 78% of the logarithmic variation of the city size.
However, as the Japan example in Section˜4.3 illustrates, seemingly innocuous choices regarding endogenous forces in QSMs can lead to markedly different counterfactual predictions, especially for distributional outcomes across regions. Most regional QSMs are Type L and rely on local dispersion forces for tractability (Redding-Dict2025, Fn. 8), which leads them to predict decentralization as transport costs fall. [^22] A subtle point is that the local–global distinction reflects detailed modeling choices rather than broad economic mechanisms. For example, a “congestion” force can be global if it arises from crowding of facilities that are accessible across regions with positive transport costs. This narrows the range of distributional outcomes such models can generate and, in turn, constrains the policy conclusions that can be drawn.
Further, some empirical contexts may be at odds with the unique-equilibrium assumption. For large transport projects such as highway systems, winners and losers may be uncertain ex-ante, making multiple equilibria potentially relevant. Evidence from the Chinese urban system indicates that urban hierarchies can be based on strong agglomeration forces or immobile factors (Baum-Snow-etal-JUE2020), both of which can generate multiple equilibria. The persistence of spatial patterns further indicates that strong agglomeration forces and path dependence shape long-run regional outcomes (Lin-Rauch-RSUE2022). A central challenge is therefore to assess whether incorporating such forces improves the empirical fit and counterfactual performance of QSMs. Relatedly, Graham-Horcher-RP2024 argue that while QSMs hold promise for applied transport policy analysis, they are not yet practice-ready, citing model validation and uncertainty quantification as key obstacles. These challenges are closely tied to how endogenous forces are specified in QSMs, since the strength and nature of agglomeration and dispersion mechanisms critically determine equilibrium spatial structure and, ultimately, the robustness of counterfactual predictions.
5  Concluding remarks
We briefly discuss several examples to illustrate that our theoretical framework also offers a unifying interpretation of seemingly heterogeneous empirical findings. See Duranton-Turner-2025 for a more comprehensive survey of empirical evidence on how transport infrastructure affects urban and regional growth.
Evidence on regional growth from Faber-REStud2014 is consistent with the predictions of Type G or GL models. He examines peripheral cities in China and finds negative effects on economic output. This may reflect a tendency for economic activity to concentrate in relatively larger or more central regions as transport access improves. [^23] Duranton-Turner-REStud2012 document that transport infrastructure in neighboring MSAs negatively affects the employment growth rate of an MSA (Table E2), a pattern consistent with global agglomeration. Baum-Snow-etal-JUE2020 provide complementary evidence for China, documenting slower growth in the hinterland prefectures compared to regional primates following the expansion of the highway system.